A company manufactures two kinds of ice-cream. The vanilla ice-cream sells for $2.50 each while the chocolate flavour ice-cream sells for $4.50 cents each. It costs the company 1 labour hour to make the vanilla flavour ice-cream and 2 labour hours to make the chocolate flavour ice-cream. The company has a total of 300 labour hours available. It costs the company 3 machine hours for the vanilla ice-cream and 2 machine hours for the chocolate ice-cream. The company has a total of 480 machine hours available. How much of each type of ice-cream should the company produce to maximise revenue? What is the maximum revenue? [Hint: let vanilla ice-cream = x]
1
Expert's answer
2020-05-05T18:44:48-0400


Maximize revenue:


Subject to:

x+2y <= 300

3x + 2y <=480

x >= 0

y >= 0



He should buy 90 vanilla icecream and 105 chocolate icecream to maximise the revenue and the maximum revenue genrated by this will be 90*2.5+105*4.5=225+472.5=697.5